Two rockets are to be launched at the same time from NASA headquarters in​ Florida, and are to meet at a space station many miles from Earth. The first rocket is to travel at 20,000 miles per hour and the second rocket will travel at 14,000 miles per hour. If the first rocket is scheduled to reach the space station 0.9 of an hour before the second​ rocket, how far is the space station from NASA​ headquarters?
Answer by Boreal(15235)

First rocket travels x hours
Second rocket travels x+0.9 hours
Distance is speed*velocity. They both go the same distance
So 20000*x=14000(x+0.9)
20000x=14000x+12600
6000x=12600
x=2.1 hours
20000*2.1=42000 miles
14000*3.0=42000 miles ANSWER